OpenStudy (anonymous):

9(4x-4)=3(12x+7) when I worked t out i got just 57 no x my choices are all real numbers or set is empty which is it ?

OpenStudy (anonymous):

okay i can help you here if you on here?

OpenStudy (anonymous):

i am here

OpenStudy (anonymous):

how do I write this

OpenStudy (anonymous):

show me what you did with this 9(4x-4)=3(12x+7)

OpenStudy (anonymous):

OK, here we go

OpenStudy (anonymous):

work with just the nine ony

OpenStudy (anonymous):

do this only 9(4x-4) I want to see your steps.

OpenStudy (anonymous):

9(4x-4)=3(12x+7) 36x-4=36x+7 36x-4+4=36x+7+4 36x-36x=11 ok here is where I get lost because there is no x so is the answer real or empty

OpenStudy (anonymous):

my bad i messed up

OpenStudy (anonymous):

9(4x-4) is 36x-36

OpenStudy (anonymous):

yes so what is x now?

OpenStudy (anonymous):

36 ?

OpenStudy (anonymous):

no .. try again hint we want X alone.

OpenStudy (anonymous):

there is no x because 36x-36x is nothing

OpenStudy (anonymous):

lost

OpenStudy (anonymous):

9*4=36 and 3*12 is 36

OpenStudy (anonymous):

okay first off there are not two 36x-36x wrong. 9(4x-4) doesn't = 36x-36x no

OpenStudy (anonymous):

36x-36 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

LOL I feel so stupid

OpenStudy (anonymous):

don't feel that way, sometimes i make little mistakes as well. we all learn in different ways.

OpenStudy (anonymous):

yes I know I am a visual learner

OpenStudy (anonymous):

yes that is okay.

OpenStudy (anonymous):

now for the other side i got 36x+21

OpenStudy (anonymous):

hey one thing is slow down and do the steps, don't skip ahead just some friendly advice

OpenStudy (campbell_st):

tootie you are correct with your solution the expansion is 36x - 36 = 36x + 21 subtract 36x from both sides and you get -36 = 21 no solution

OpenStudy (anonymous):

yes that is 36x+21 now again how do we get x alone?

OpenStudy (anonymous):

subtract 36x-36x

OpenStudy (campbell_st):

its basically saying y = 36x - 36 and y = 36x + 21 the lines are parallel ... so no intersection....

OpenStudy (anonymous):

campbell yes are right about 36x-36x I wasn't at that point yet..

OpenStudy (campbell_st):

well tootie has actually has a correct piece of maths above... sorry...

OpenStudy (anonymous):

so the solution set is empty. right?

OpenStudy (anonymous):

I was walking through the steps campbell

OpenStudy (campbell_st):

correct... just practise expanding terms in brackets...

OpenStudy (campbell_st):

the expansion of the pronumeral was correct... its just the constant term you need to be aware of

OpenStudy (anonymous):

thank you for taking me through the steps I was not sure if I worked it right Godorovg

OpenStudy (anonymous):

tootie 149 I was going through the steps with you, and to make sure you understood what was going

OpenStudy (anonymous):

empty is the answer

OpenStudy (anonymous):

Thank you so much I needed to know I worked it right, and you did that thank you again

OpenStudy (campbell_st):

hope it helps...

OpenStudy (anonymous):

oh no trouble at all..

OpenStudy (tiffanymak1996):

in this case if there is no x, then the set is empty, if it's all real number, all real number would work with the equation.

OpenStudy (anonymous):

right back at you I am a fan of yours now too. :)

OpenStudy (anonymous):

tiffanymak could you give me an example of what you mean when it comes to real numbers

OpenStudy (anonymous):

real numbers are 1,2,3 ,5 and so on

OpenStudy (anonymous):

still lost when it comes to real numbers.

OpenStudy (tiffanymak1996):

real numbers are like all kinds of number, rational, irrational, integer etc. just not complex number.

OpenStudy (tiffanymak1996):

say: integer: - [infinite], -8, -7, -6,-5...0,1,2,3,4,5....., [infinite] these are real no.s

OpenStudy (anonymous):

hey tootie look at this http://math10.edcentre.ca/bridgemath/numbers/setnos.php

OpenStudy (tiffanymak1996):

sqrt 2, sqrt3, pi etc, are real numbers

OpenStudy (tiffanymak1996):

yeah, and well that link too.

OpenStudy (tiffanymak1996):

almost every kind of no.s are real, except for i, (which is sqrt(-1)) the imaginary number

OpenStudy (anonymous):

Tootie what is the biggest problem in the class you are in??

OpenStudy (anonymous):

there is so much I do not know where to start

OpenStudy (anonymous):

okay I am guessing this college algbra right?

OpenStudy (anonymous):

YOU got it

OpenStudy (anonymous):

okay first off in this class I can find you videos where can get the basics by watching would that help you?

OpenStudy (anonymous):

but the said part my daughter was doing this in the 6th grade here, and they make little remarks about it

OpenStudy (anonymous):

yes that would be great!!!!!!!

OpenStudy (anonymous):

Tootie here is a shock I taught myself math through videos on the internet.

OpenStudy (anonymous):

really please give me the links.

OpenStudy (tiffanymak1996):

a really good site would be the paul's notes

OpenStudy (anonymous):

okay will back

OpenStudy (anonymous):

paul's notes

OpenStudy (anonymous):

I will be waiting Godorovg

OpenStudy (campbell_st):

tootie... have a look at Khan academy on youtube..or google Khan Academy... they have lots of videos explaining concepts

OpenStudy (anonymous):

never heard of them

OpenStudy (tiffanymak1996):

MIT has free online courses for tons of subject, i don't know about algebra but you can check out their site.

OpenStudy (anonymous):

just looked it up thanks so much campbell_st

OpenStudy (anonymous):

the MIT?

OpenStudy (tiffanymak1996):

yeah

OpenStudy (anonymous):

ok will look it up thanks so much, so much

OpenStudy (anonymous):

tootie watch this video http://www.youtube.com/watch?v=1Amt_-uB9QQ

OpenStudy (tiffanymak1996):

http://ocw.mit.edu/courses/

OpenStudy (tiffanymak1996):

that's the link

OpenStudy (anonymous):

looking now will come back after i look at it

OpenStudy (anonymous):

tootie it is the first part of this class

OpenStudy (anonymous):

thanks to everyone for their help. really thanks

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Always glad to help out :)

OpenStudy (tiffanymak1996):

so, do you home study math?

OpenStudy (anonymous):

yes and I am working on Physics and chem through MIT open course study right now. Plus Calc 1 and Cal2 as well

OpenStudy (tiffanymak1996):

cool, i study through stanford.

OpenStudy (anonymous):

I wish I could get in there.

OpenStudy (anonymous):

can't now because I ran away to college when I wasn't ready GPA way too low.

OpenStudy (tiffanymak1996):

me too, i'm only joining their for gifted child program, but still has two years to go till college.

OpenStudy (anonymous):

yeah, funny thing is I didn't how I was gifted math either.

OpenStudy (tiffanymak1996):

what's the difference between college and university?

OpenStudy (anonymous):

Tiffany when I first took I couldn't get it and than day by day it started to make sense until now here I am.

OpenStudy (anonymous):

nothing really expect for name . Now there two year and four colleges and University out there. It all depends on what you want to do in college?

OpenStudy (anonymous):

also, depending on your major you could go on-line and get your degree.

OpenStudy (tiffanymak1996):

i can't decide yet, math and literature is my life, but my parents wants me to be a lawyer, just to make sure i'll have enough to live by

OpenStudy (anonymous):

Lawyers would be okay, only thing is you go to school for eight to nine years than have to past a test to become a lawyer.

OpenStudy (tiffanymak1996):

i'll be so old by then :)

OpenStudy (anonymous):

nah you are still pretty young I am guessing

OpenStudy (tiffanymak1996):

nah, just old for starting to work

OpenStudy (anonymous):

yeah it happens no big deal. I think most jobs want older people now a days

OpenStudy (anonymous):

hey tootie what you think about the video?